# FlexiSCD: Flexible Surface Code Deformer for Dynamic Defects

### v2

@@ 24            PatchQ_RM $p$, $e$
Comparison with ASC. ASC solely employs the DataQ_RM instruction even for defective syndrome qubits. For removing a defective syndrome qubit, ASC must apply DataQ_RM four times to remove adjacent data qubits, even if they are not defective. This can result in an unnecessary loss of code distance. For example, the deformed code produced by ASC in Fig. 7(a) has a $Z$-distance of only 3, while our strategy can preserve it to be 5.
Figure 7: Illustration of ASC and Q3DE. (a) The resulting code of ASC has $X$- and $Z$-distance both equal to $3$, while our instruction Fig. 6(b) leads to a code with $Z$-distance $5$ and $X$-distance $3$. (b) Code enlargement of Q3DE equates to multiple PatchQ_ADD instructions.
@@
Comparison with ASC. ASC  disables defective qubits on the boundary, disregarding their impact on the $X$-, $Z$-distances. As Fig. 8 shows, ASC converts the gauge operator $Z_{0}$ into stabilizer to minimize the number of disabled qubits, leading to a smaller $Z$-code distance than FlexiSCD.
Figure 8: Convert the right gauge operator into stabilizer to balance the $X$-, $Z$-distance. (a) ASC converts the gauge operator $Z_{0}$ into stabilizer, leading to unbalanced $X$-, $Z$-distance and a smaller code distance. (b) FlexiSCD converts the gauge operator $X_{0}$ into stabilizer, leading to balanced $X$-, $Z$-distance and a larger code distance.
@@ V-B Adaptive Enlargement Subroutine
This subroutine receives the deformed code of the Defect Removal Subroutine and efficiently restores the lost code distance by introducing new qubits. As illustrated in Algorithm 2, this enlargement process addresses defects in the newly introduced qubits and ensure a restoration to the original code distance.
@@ 27             Apply Algorithm 1 to remove $d_{i}$
Code enlargement with defective patches. After defect removal, we evaluate the reduction in both $X$- and $Z$-code distances compared to the ideal surface code. Calculating the code distance for every patch added to the surface code could be inefficient. In the find_layer function, we first identify a set of patches that can increase the code distance by at least one unit of the relevant type, assuming that the identified layer is defect-free.
However, new layers may contain defects, which invalidates our assumption. For example, in Fig. 10 (a), defective qubits $q_{1}$ and $q_{2}$ are removed, resulting in an irregular boundary that hinders regular enlargement. Similarly, in Fig. 10 (c), the defective qubit $q_{0}$ in the prospective scale layer also complicates the enlargement.
To solve this problem, FlexiSCD  combines the instruction PatchQ_ADD with the qubit removal subroutine DataQ_RM, SyndromeQ_RM, PatchQ_RM, as shown in line 24 in Algorithm 2. Initially, defective qubits are temporarily disregarded to facilitate the regular enlargement. Subsequently, these defective qubits are excluded using our qubit removal instructions, resulting in an enlarged code with defects removed as the output of the Code Deformation Unit (Fig. 10 (b) (d)). To execute the deformation process, qubits are introduced in the appropriate state all at once, and stabilizers in the new code are measured. Importantly, this seamless integration of qubit removal and enlargement is facilitated by the unified gauge transformation perspective underlying these deformation instructions.
Figure 10: Enlargement while addressing defects. (a) Removed defects create an irregular boundary. (b) Enlarging on the irregular boundary in (a). (c) Defect can occur on prospective enlarging layers. (d) Enlarging by a qubit layer with a defect in (c).
@@ VI The Layout Generator
FlexiSCD’s layout generator aims to facilitate communication between surface codes in the presence of potential code enlargement due to defect mitigation. The key idea is to introduce an additional $\Delta d$ inter-space, as shown in Fig. 9(a), to accommodate enlargements of up to size $\Delta d$ (or even $2\Delta d$ if enlarging in both directions). The parameter $\Delta d$ is determined by the defect error model, chosen such that enlarging within size $\Delta d$ can effectively restore the code distance of the deformed code with high probability. Therefore, the communication channel of width $d$ will only be compromised by code enlargement with a negligible probability.
Comparison with Q3DE. As illustrated in Fig. 7 (b), Q3DE supports fixed enlargement only, whereas our FlexiSCD allows the introduction of only the necessary number of patches to restore the code distance, resulting in savings in qubit resources needed for incorporation into the code. Moreover, the basic process in Q3DE lacks the capability to address irregular boundaries caused by defects or defects occurring on newly introduced qubits. Consequently, Q3DE may inadvertently introduce additional defective qubits during the enlargement process, leading to an increase in error sources that counteract the intended function of the introduced redundant qubits. This failure to effectively mitigate defects can result in a logical error rate that exceeds expectations.
FlexiSCD’s layout generator. The output layout is determined by three parameters:
(1) Logical qubit number $N$. The logical qubits in the layout should be the algorithmic qubits needed for the quantum program, plus the additional qubits essential for facilitating quantum computation, such as logical qubits prepared in magic states for implementing logical $T$ gates [41]. These logical qubits are arranged in a lattice on a plane, with each one encoded within a surface code and equally spaced apart from one another.
@@
TABLE II: Results of Q3DE , ASC and FlexiSCD for various programs with different scale
Benchmark Q3DE ASC FlexiSCD
@@
$p_{block}=1-\sum_{k=0}^{\lfloor\Delta d/D\rfloor}p(k)$ (1)
(1) Space-time overhead of logical operations. The blocked channel may lead to space overhead for long-distance logical operations. For example, in Fig. 9(b), the CNOT gate from qubit $B$ to $C$ must traverse a longer path (solid line) instead of the shorter one (dotted lines) due to the enlarged qubit $B$. Moreover, the blocked channel can result in severe time overhead. Logical operations may become unfeasible due to blocked paths, such as the CNOT gate from qubit $A$ to $B$. In such cases, the program must be paused for a long time (thousands of QEC cycles) until the effects of defects subside, and the code shrinks back to its original size, making the ancilla path available again. In contrast, our layout with an extra inter-space $\Delta d$ effectively suppresses the block probability, thus reducing the space-time cost of logical operations to a nearly optimal level.
VII Implementation and Evaluation
VII-A Experimental Setup
Evaluation setting. We use the Stim stabilizer circuit simulator [51] and the PyMatching [52] decoder for our numerical simulations. For program compilation, we employ the lattice surgery framework described in [53, 42], along with the $T$ factory for implementing logical $T$ gates using magic state distillation strategies [40].
Benchmark programs. We have selected several quantum programs designed to tackle real-world challenges. These programs include Simon’s algorithm [54], Ripple Carry Adder (RCA) [55], Quantum Fourier Transform (QFT) [56], and Grover’s algorithm [5]. The first suffix of the name indicates the number of logical qubits used, while the second suffix represents the number of repetitions. For instance, ”QFT-100-20” indicates the task of 20 QFT layers on 100 qubits, a common scenario used in more complex algorithms like Shor’s algorithm [4].
Figure 11: (a) Logical error rate of with qubit removal by FlexiSCD and without any treatment. (b) Cost in distance recovery. (c) Throughput vs Defect.
@@ VII-A Experimental Setup
Error model. We adopt an error model similar to [2]. The physical error rate of each qubit is assumed to be $p=10^{-3}$. For dynamic defects, each physical qubit experiences a dynamic defect event following an exponential distribution with an average event rate of $\lambda=1/(26\times 10\text{ s})$ [22]. When a physical qubit is attacked, the adjacent 24 qubits are affected, resulting in a defective region of size 4. The physical error rate of qubits in the affected region increases to approximately 50% and lasts for $T=25$ ms, equivalent to approximately 25000 QEC cycles.
@@ VII-B Overall Performance
We compare our FlexiSCD with two previous frameworks, Q3DE method [2] and adaptive surface code (ASC) [1], on benchmark programs mentioned in Section VII-A. The results are presented in Table II. We evaluate these tasks on surface codes of appropriate distances chosen to meet the target retry risk levels of $1\%$ and $0.1\%$ in our framework. Overall, our FlexiSCD in Table II demonstrates a consistently better performance. In particular, we highlight the following three observations:
@@
    All tasks processed by Q3DE exhibit a retry risk of $\sim 100\%$, indicating an inability to mitigate the impact of dynamic defects. This limitation stems from its enlargement operation with a fixed layout, which causes the ancilla space near the defective qubits to become blocked. Additionally, the large scale of tasks necessitates high-fidelity logical qubits, implying a large code distance $d$ for the surface code. With more qubits used to encode a logical qubit, there is a higher probability of the qubit blocks being attacked by defect events. Consequently, most qubit blocks become enlarged in Q3DE , and a significant portion of the ancilla paths are blocked. Fewer or even no available ancilla paths for executing non-local operations result in the high retry risk.
@@
    Without the ability to recover the decreased distance caused by removing defective qubits, the retry risk of tasks executed by ASC is 35x-70x higher than those by our FlexiSCD. This disparity arises from the task’s retry risk being determined by the worst qubit blocks during runtime, since a single logical error on any logical qubit can lead to the failure of the entire task. Failure to restore a low-distance logical qubit to the original required distance $d$ for the task significantly amplifies the retry risk.
@@
    With FlexiSCD’s layout, it requires only around $20\%$ more physical qubits to resolve the communication blocking issue and achieve a low retry risk. Moreover, even with the same number of physical qubits, FlexiSCD still outperforms ASC. For example, in the task ”QFT-100-20,” FlexiSCD with $d=25$ and $\Delta d=4$ and ASC with $d=27$ both use roughly $9\times 10^{5}$ physical qubits. However, the retry risk of FlexiSCD is still 6.3x lower than that of ASC.
@@ VII-C Analysis on Sub-components
In this section, we demonstrate the superior performance of our FlexiSCD due to its adaptive deformation framework and layout. Our approach excels in saving qubit resources, maintaining the QEC capability of the deformed code, and enhancing surface code communication compared to other methods. We have conducted a series of experiments to showcase these advantages.
@@
Qubit resource saving. We demonstrate the substantial resource savings achieved by FlexiSCD by comparing the physical qubit count required for the layout to achieve a $1\%$ retry risk with different methods. Our comparison includes Lattice Surgery, the revised version of Q3DE, and ASC. In the revised Q3DE, we adjust the inter-space to $2d$ to prevent the enlargement from blocking surrounding communication paths, as illustrated in Fig. 9(c). As depicted in Fig. 12, FlexiSCD requires significantly fewer qubits compared to all previous methods, approximately 75% less than Lattice Surgery, 50% less than Q3DE, and 15% less than ASC.
Improved logical error rate by qubit removal. In Fig. 11(a), we compare the logical error rates of surface codes with and without defective qubits removed by FlexiSCD, across different numbers of defective qubits. We do not present the cases for $d=21$ and $d=27$ because the logical error rates are so low that numerical simulations cannot provide reasonable estimations.
@@
The results show that FlexiSCDsignificantly enhances the logical error rate by removing defects. Codes with defects removed achieve comparable logical error rates to those without any treatment, despite having much smaller code distances. For instance, when there are 10 defective qubits, a distance-9 code with defects removed shows a similar logical error rate to a distance-15 code without any treatment. This also underscores that simply expanding the code size while retaining defective qubits, as in Q3DE, cannot effectively improve logical error rates as expected.
@@
Enhanced code distance by adaptive defect removal. We compare the cost of distance recovery of FlexiSCD and previous methods such as ASC that use oversimplified deformation operations. Fig. 11(b) gives the code distance of defective codes of different original sizes, w.r.t. varying number of defective qubits. It indicates that our method can preserve the code distance better than the previous methods, leading to a lower cost when recovering the code distance through enlargement. Additionally, the advantage of our method becomes more significant on codes of larger size and when there are more defective qubits.
@@
Parallelism improvement by FlexiSCD’s layout. In Fig. 11(c), we compare the efficiency of qubit communication using the layouts in three schemes: (1) FlexiSCD, (2) Q3DE with $d$ inter-space, and (3) classical lattice surgery (LS), representing optimal runtime. We consider $100$ logical qubits and three sets of tasks, each comprising $5$ tasks with $25$ CNOT gates on $50$ distinct logical qubits. The tasks in each set can be completed in $16$, $19$, and $22$ time-steps on the LS layout, respectively, indicating three levels of parallelism. We then sample defects $100$ times for each task and compute the average throughput of each task set using the layouts in FlexiSCD and Q3DE.
@@
The results indicate that as the defect rate increases, the throughput significantly drops for the Q3DE layout, whereas it only experiences a slight decline with our layout. This is because expanding all the defective code to a $2d$ size severely blocks the ancilla paths for long-range CNOT gates, resulting in a much longer execution time, while the additional $\Delta d$ space in our layout significantly reducing the probability of ancilla paths being blocked. Therefore, the FlexiSCD layout effectively facilitates the computation among logical qubits.

### v3

@@ 24            PatchQ_RM $p$, $e$
V The Code Deformation Unit
The code deformation unit is the core component of Surf-Deformer for adaptive defect mitigation. Before every QEC cycle, the code deformation unit receives the current surface code configuration (including the shape and location of logical qubit patches) from the last round of deformation and current defect information from the dynamic defect detector. Leveraging this information, the code deformation unit executes two subroutines sequentially, as illustrated in fig. 5. The first subroutine, Defect Removal, removes detected qubits from the current code. The second subroutine Adaptive Enlargement takes the resulting code, calculates the new code distance, and adaptively enlarges the code.
V-A Defect Removal Subroutine
@@
Comparison with ASC-S. ASC-S solely employs the DataQ_RM instruction even for defective syndrome qubits. For removing a defective syndrome qubit, ASC-S must apply DataQ_RM four times to remove adjacent data qubits, even if they are not defective. This can result in an unnecessary loss of code distance. For example, the deformed code produced by ASC-S in fig. 7(a) has a $Z$-distance of only 3, while our strategy can preserve it to be 5.
Figure 7: Illustration of ASC-S and Q3DE. (a) The resulting code of ASC-S has $X$- and $Z$-distance both equal to $3$, while our instruction fig. 6(b) leads to a code with $Z$-distance $5$ and $X$-distance $3$. (b) Code enlargement of Q3DE equates to multiple PatchQ_ADD instructions.
@@
Comparison with ASC-S. ASC-S  disables defective qubits on the boundary, disregarding their impact on the $X$-, $Z$-distances. As fig. 8 shows, ASC-S converts the gauge operator $Z_{0}$ into stabilizer to minimize the number of disabled qubits, leading to a smaller $Z$-code distance than Surf-Deformer.
Figure 8: Convert the right gauge operator into stabilizer to balance the $X$-, $Z$-distance. (a) ASC-S converts the gauge operator $Z_{0}$ into stabilizer, leading to unbalanced $X$-, $Z$-distance and a smaller code distance. (b) Surf-Deformer converts the gauge operator $X_{0}$ into stabilizer, leading to balanced $X$-, $Z$-distance and a larger code distance.
@@ V-B Adaptive Enlargement Subroutine
This subroutine, outlined in algorithm 2, receives the deformed code of the Defect Removal Subroutine and efficiently restores the lost code distance by introducing new qubits.
@@ 27             Apply algorithm 1 to remove $d_{i}$
Code enlargement with defective patches. After defect removal, we evaluate the reduction in both $X$- and $Z$-code distances compared to the ideal surface code. Calculating the code distance for every patch added to the surface code could be inefficient. In the find_layer function, we identify a set of patches, denoted as a scale layer, that can increase the code distance by at least one unit of the relevant type.
However, both boundary and new scale layers may contain defects, which requires adaptive enlargement. For example, in fig. 9(a), defective qubits $q_{1}$ and $q_{2}$ are removed, resulting in an irregular boundary that hinders regular enlargement. Similarly, in fig. 9(c)(d), the defective qubit $q_{0}$ in the prospective scale layer also complicates the enlargement, since it need two scale layers to restore its $Z_{L}$ distance of 5.
To solve this problem, Surf-Deformer combines the instruction PatchQ_ADD with the qubit removal subroutine DataQ_RM, SyndromeQ_RM, PatchQ_RM, as shown in line 24 in algorithm 2. Initially, defective qubits are temporarily disregarded to facilitate the regular enlargement. Subsequently, these defective qubits are excluded using our qubit removal instructions, resulting in an enlarged code with defects removed as the output of the Code Deformation Unit (fig. 9(b)(d)). To execute the deformation process, qubits are introduced in the appropriate state all at once, and stabilizers in the new code are measured. Importantly, this seamless integration of qubit removal and enlargement is facilitated by the unified gauge transformation perspective underlying these deformation instructions.
Figure 9: Enlargement while addressing defects. (a) Removed defects create an irregular boundary. (b) Enlarging on the irregular boundary. (c) Defect can occur on prospective scale layers. (d) Enlarging by two qubit layers.
Comparison with Q3DE. As shown in fig. 7(b), Q3DE supports fixed enlargement only, whereas our Surf-Deformer allows the introduction of only the necessary number of patches to restore the code distance, resulting in savings in qubit resources needed for incorporation into the code. Moreover, the basic process in Q3DE lacks the capability to address irregular boundaries caused by defects or defects occurring on newly introduced qubits. Consequently, Q3DE may inadvertently introduce additional defective qubits during the enlargement process, leading to an increase in error sources that counteract the intended function of the introduced redundant qubits. This failure to effectively mitigate defects can result in a logical error rate that exceeds expectations.
@@ VI The Layout Generator
Surf-Deformer’s layout generator aims to facilitate communication between surface codes in the presence of potential code enlargement due to defect mitigation. The key idea is to introduce an additional $\Delta d$ inter-space, as shown in fig. 10(a), to accommodate enlargements of up to size $\Delta d$ (or even $2\Delta d$ if enlarging in both directions). The parameter $\Delta d$ is determined by the defect error model, chosen such that enlarging within size $\Delta d$ can effectively restore the code distance of the deformed code with high probability. Therefore, the communication channel of width $d$ will only be compromised by code enlargement with a negligible probability.
Surf-Deformer’s layout generator. The output layout is determined by three parameters:
(1) Logical qubit number $N$. The logical qubits in the layout consist algorithmic qubits needed for the quantum program, plus the additional qubits essential for facilitating quantum computation (e.g. logical qubits prepared in magic states for implementing logical $T$ gates [41]). These logical qubits are arranged in a lattice on a plane, with each one encoded within a surface code and equally spaced apart from one another.
@@
$p_{block}=1-\sum_{k=0}^{\lfloor\Delta d/D\rfloor}p(k)$ (1)
(1) Space-time overhead of logical operations. The blocked channel may lead to space overhead for long-distance logical operations. For example, in fig. 10(b), the CNOT gate from qubit $B$ to $C$ must traverse a longer path (solid line) instead of the shorter one (dotted lines) due to the enlarged qubit $B$. Moreover, the blocked channel can result in severe time overhead. Logical operations may become unfeasible due to blocked paths, such as the CNOT gate from qubit $A$ to $B$. In such cases, the program must be paused for a long time (thousands of QEC cycles) until the effects of defects subside and the code shrinks back to its original size, making the ancilla path available again. In contrast, our layout with an extra inter-space $\Delta d$ effectively suppresses the block probability, thus reducing the space-time cost of logical operations to a nearly optimal level.
VII Implementation and Evaluation
VII-A Experimental Setup
Evaluation setting. We use the Stim [61] stabilizer circuit simulator and the PyMatching [62] decoder for our numerical simulations. For program compilation, we employ the lattice surgery framework described in [63, 42], along with the $T$ factory for implementing logical $T$ gates using magic state distillation strategies [40].
Benchmark programs. We have selected several quantum programs designed to tackle real-world challenges. These programs include Simon’s algorithm [64], Ripple Carry Adder (RCA) [65], Quantum Fourier Transform (QFT) [66], and Grover’s algorithm [5]. The first suffix of the name indicates the number of logical qubits used, while the second suffix represents the number of repetitions. For instance, ”QFT-100-20” indicates the task of 20 QFT layers on 100 qubits, a common scenario used in Shor’s algorithm [4].
TABLE II: Results of Q3DE , ASC-S and Surf-Deformer for various programs with different scale
Benchmark Q3DE ASC-S Surf-Deformer
@@ VII-A Experimental Setup
Physical error model. We employ a widely-used circuit-level error model [15, 61]. We associate same probabilities $p=10^{-3}$, which is 0.1 times the surface code’s error threshold, with the single-qubit depolarizing error channel for single-qubit gates, the two-qubit depolarizing error channel for two-qubit gates, and the Pauli-X error channel for measurement and reset operations.
Dynamic defect model. We adopt the dynamic defect model in [2], which is derived from physical experiments [22]. Each physical qubit experiences a dynamic defect event following an exponential distribution with an average event rate of $\lambda=1/(26\times 10\text{ s})$ [22]. When a physical qubit is attacked, the adjacent 24 qubits are affected, resulting in a defective region of size 4. The physical error rate of qubits in the affected region increases to approximately 50% and lasts for $T=25$ ms, equivalent to approximately 25,000 QEC cycles [17].
@@ VII-B Overall Performance
We compare our Surf-Deformer with two previous frameworks, Q3DE method [2] and adaptive surface code (ASC-S)[33], on benchmark programs mentioned in section VII-A. The results are presented in table II. We evaluate these tasks on surface codes of appropriate distances chosen to meet the target retry risk levels of $1\%$ and $0.1\%$ in our framework. Overall, our Surf-Deformer in table II demonstrates a consistently better performance. In particular, we highlight the following three observations:
@@
    All tasks processed by Q3DE exhibit a retry risk of $\sim 100\%$, failing to mitigate the impact of dynamic defects. This limitation stems from its enlargement operation with a fixed layout, which causes the ancilla space near the defective qubits to become blocked. Additionally, the large scale of tasks necessitates high-fidelity logical qubits, implying a large code distance $d$ for the surface code. With more qubits used to encode a logical qubit, there is a higher probability of the qubit blocks being attacked by defect events. Consequently, most qubit blocks become enlarged in Q3DE, and a significant portion of the ancilla paths are blocked. Fewer or even no available ancilla paths for executing non-local operations result in the high retry risk.
@@
    Without the ability to recover the decreased distance caused by removing defective qubits, the retry risk of tasks executed by ASC-S is 35x-70x higher than those by our Surf-Deformer. This disparity arises from the task’s retry risk being determined by the worst qubit blocks during runtime, since a single logical error on any logical qubit can lead to the failure of the entire task. Failure to restore a low-distance logical qubit to the original required distance $d$ for the task significantly amplifies the retry risk.
@@
    With Surf-Deformer’s layout, it requires only around $20\%$ more physical qubits to resolve the communication blocking issue and achieve a low retry risk. Moreover, even with the same number of physical qubits, Surf-Deformer still outperforms ASC-S. For example, in the task ”QFT-100-20,” Surf-Deformer with $d=25$ and $\Delta d=4$ and ASC-S with $d=27$ both use roughly $9\times 10^{5}$ physical qubits. However, the retry risk of Surf-Deformer is still 6.3x lower than that of ASC-S.
Figure 11: (a) Logical error rates for the surface code with defect removal compared to no treatment. (b) Loss in code distance during adaptive defect removal in Surf-Deformer compared to ASC-S. (c) Throughput of quantum tasks with varying parallelism on the adaptive layout of Surf-Deformer compared to Q3DE.
@@ VII-C Analysis on Sub-components
In this section, we demonstrate the superior performance of our Surf-Deformer due to its adaptive deformation framework and layout. Our approach excels in saving qubit resources, maintaining the QEC capability of the deformed code, and enhancing surface code communication compared to other methods. We have conducted a series of experiments to showcase these advantages.
@@
Qubit resource saving. We demonstrate the substantial resource savings achieved by Surf-Deformer by comparing the physical qubit count required for the layout to achieve a $1\%$ retry risk with different methods. Our comparison includes Lattice Surgery, the revised version of Q3DE, and ASC-S. In the revised Q3DE, we adjust the inter-space to $2d$ to prevent the enlargement from blocking surrounding communication paths, as illustrated in fig. 10(c). As depicted in fig. 12, Surf-Deformer requires significantly fewer qubits compared to all previous methods, approximately 75% less than Lattice Surgery, 50% less than Q3DE, and 15% less than ASC-S.
@@
Improved logical error rate by qubit removal. In fig. 11(a), we compare the logical error rates of surface codes with and without defective qubits removed by Surf-Deformer, across different numbers of defective qubits. We didn’t present the cases for $d=21$ and $d=27$ because the logical error rates are so low that numerical simulations cannot provide reasonable estimations. The results show that codes with defects removed by Surf-Deformer achieve comparable logical error rates to those without any treatment, despite having much smaller code distances. For instance, when there are 10 defective qubits, a distance-9 code with defects removed shows a similar logical error rate to a distance-15 code without any treatment. This also underscores that simply expanding the code size while retaining defective qubits, as in Q3DE, cannot effectively improve logical error rates as expected.
@@
Enhanced code distance by adaptive defect removal. We compare the cost of distance recovery of Surf-Deformer and previous methods such as ASC-S that use oversimplified deformation operations. fig. 11(b) gives the code distance of defective codes of different original sizes, w.r.t. varying number of defective qubits. It indicates that our method can preserve the code distance better than the previous methods, leading to a lower cost when recovering the code distance through enlargement. Additionally, the advantage of our method becomes more significant on codes of larger size and when there are more defective qubits.
@@
Parallelism improvement by Surf-Deformer’s layout. In fig. 11(c), we compare the efficiency of qubit communication using the layouts in three schemes: (1) Surf-Deformer, (2) Q3DE with $d$ inter-space, and (3) classical lattice surgery (LS), representing optimal runtime. We consider $100$ logical qubits and three sets of tasks, each comprising $5$ tasks with $25$ CNOT gates on $50$ distinct logical qubits. The tasks in each set can be completed in $16$, $19$, and $22$ time-steps on the LS layout, respectively, indicating three levels of parallelism. We then sample defects $100$ times for each task and compute the average throughput of each task set using the layouts in Surf-Deformer and Q3DE.
@@
The results indicate that as the defect rate increases, the throughput significantly drops for the Q3DE layout, whereas it only experiences a slight decline with our layout. This is because expanding all the defective code to a $2d$ size severely blocks the ancilla paths for long-range CNOT gates, resulting in a much longer execution time, while the additional $\Delta d$ space in our layout significantly reducing the probability of ancilla paths being blocked. Therefore, the Surf-Deformer layout effectively facilitates the computation among logical qubits.
VII-D Trade-off between Qubit Resource and QEC Performance
In fig. 13(a), we compare the trade-off lines of physical qubit count and retry risk for Surf-Deformer and ASC-S. Our findings indicate that the logical error rate of logical qubits transformed by Surf-Deformer maintains the trade-off line of the non-defective surface code’s QEC performance, with the logical error rate decreasing exponentially with increasing code distance. Furthermore, Surf-Deformer demonstrates a more efficient trade-off line than ASC-S, achieving the same retry risk with lower physical qubit overhead.
Surf-Deformer’s deformation method can also be integrated with post-selection method in [33] to enhance the removal of static defects and post-selection of chiplets. As shown in fig. 13(b), Surf-Deformer’s deformation method increases the yield rate of the targeted distance-27 surface code, with more significant improvements as the size of the original patch increases. For instance, when using size-25 patches with 20 faulty qubits, the yield rate of Surf-Deformer (0.75) is approximately double that of the original ASC-S (0.39).
Figure 13: (a) Trade-off between QEC performance (Retry risk) and resource efficiency (# Physical qubit) (b) Yield rates of ASC-S and Surf-Deformer for the task of deforming a $l=35$ patch with static faulty qubits to a surface code with a distance of no less than 27.
VII-E Robust to Various Error Models and Unreliable Detection
In this section, we evaluate Surf-Deformer’s robustness under high correlated error rates and unreliable defect detection. In fig. 14(a), we simulate and compare the logical error rates of the surface code and Surf-Deformer’s deformed code under various 2-qubit gate correlated error settings, with single-qubit error rates fixed at $p=0.001$. The results show that Surf-Deformer maintains a 10x improvement over the original surface code as the correlated error rate increases. In fig. 14(b), we use an existing detection method [2] which can achieve the false-positive and true-negative detection probabilities being below 0.01. The green line, representing the unreliable detection scenario, is close to the orange line, representing the precise detection scenario, indicating that Surf-Deformer’s QEC performance remains robust even with unreliable defect detection.
Figure 14: Logical error rate of a distance-9 surface code under (a) various 2-qubit gates correlated errors and (b) the presence of imprecise defect detection.